angles w and x are supplementary if m and w is 37 degrees what is m and x?
Answer by stanbon(57234)

angles w and x are supplementary if m and w is 37 degrees what is m and x?
--------------------
w + x = 180
m + w = 37
---------------------
Subtract to get:
x-m = 143
x = 143+m
m = x-143
=====================
Cheers,
Stan H.